Design a spring for suspension system of a motorcycle with a weight of two people riding the bike. You should assume the required parameters as per the following table,


Sr. No.


Reg No


Starting d


Difference for selection of diameter


Ends of spring



1


1-10


0.01 x R


0.001 x R


Squared or Closed



2


11-20


0.01 x R


0.001 x R


Squared or Closed



3


21-30


(0.01 x R)/2


0.001 x R


Squared and Ground



4


31-40


(0.01 x R)/2


0.001 x R


Squared and Ground



5


41-50


0.01 x R


0.001 x R


Squared or Closed



6


51-60


0.01 x R


0.001 x R


Squared or Closed



7


61-70


(0.001 x R)/2


0.001 x R


Squared and Ground



8


71-80


(0.001 x R)/2


0.001 x R


Squared and Ground




* where R = Your Registration Number


Start your calculation first by using material 302 stainless steel and if the required result is not output, then change the material till you find one. Select the spring as wound.



HINT: Make an excel sheet for making the table of spring selection.



1
Expert's answer
2021-11-22T17:18:03-0500

Assume number of turns = 15


assume deflection = 2 inches = 50mm


Load = 150kg = 1500N


spring stiffness , k = 1500/0.05 = 30000 N/m = 30 kN/m


Material = 302 stainless Steel


E = 210 Gpa


starting (d) , spring wire dia = 0.05mm


D is mean coil diameter


This excel sheet gives us; d and D


use formula,


k = d4G/8D3N




Note- The color coded dimesions of d & D(in mm) are suitable. U can choose any of these dimensions.


You can also see the formula used in formula bar.